OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (anonymous):

y = 1/3x y = -1/3x + 4 y = 3x + 20 y = -3x - 16

OpenStudy (ash2326):

@ireallyjustdontknow Did you try this?

OpenStudy (anonymous):

honestly, no.

OpenStudy (ash2326):

Try this, it's easy. I think you asked the same type of question earlier. Try once, if you don't get it. I'll help you

OpenStudy (anonymous):

I got it. lol

OpenStudy (ash2326):

:DWhat's the correct option?